The parallel combination of five cells, each with electromotive force of 1.5 V and internal resistance of 0.5 Ω, is connected to an external resistance of 5.0 Ω. Find the (a) total resistance of the circuit, (b) electromotive force of the battery, (c) total resistance of the circuit, and (d) current through the external circuit.



Given: n=5


E= 1.5 V


r= 0.5 Ω


R= 5.0 Ω


1
Expert's answer
2022-04-21T13:00:55-0400

(a)

"R_{tot}=\frac{r}{5}+R=\frac{0.5}{5}+5.0=5.1\:\Omega"

(b)

"{\cal E}=1.5\:\rm V"

(c)

"I=\frac{\cal E}{R_{tot}}=\frac{1.5}{5.1}=0.29\:\rm A"